An investor purchases a 10yr, 6% coupon bond at a price of 1000. compute the two year holding period yield assuming interest rates increase to 8% immediately after the bond is purchased.

A. .5%

B. 2.3%

C. 8.7%

D. 10.4%
Answer #1

Coupon =6%*1000 =60
Price of bond now =PV of Coupons+PV of Par Value =60*((1-(1+8%)^-8)/8%+1000/(1+8%)^8 =885.0672

The two year holding period return =(885.067+2*Coupon -1000)/1000 =(885.067+2*60 -1000)/1000=0.5%
(Option a is correct option)